OpenStudy (anonymous):

Find x if the slope of a line passing through points A (-2, x) and B (4, 1) equals -2/3

myininaya (myininaya):

(1-x)/(4-(-2))=-2/3 (1-x)/(6)=-2/3 1-x=-2(6)/3 -x=-12/3-1 -x=-4-1 -x=-5 x=5